OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 −2 −3 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system

OpenStudy (anonymous):

have you tried substitution?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Want to give it a shot? First thing you can do is isolate x on each equation. That means x has to be by itself on each equation. Can you do that?

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

3x − 6y = 3

OpenStudy (anonymous):

divide X on both sides?

OpenStudy (anonymous):

multiply the top by 5 and the bottom by 6 this means you ca eliminate y i'll show you

OpenStudy (anonymous):

42x-30y=-66 ; this is x6 15x-30y=15; this is x5 now subtract to get 27x=-81 now divide to get x=-81/27 = -3